A circular loop of radius R = 15 cm is centered at the origin in the region of a uniform, constant electric field. When the normal to the loop points in the positive x-direction the flux through the loop is Φ1 = 65 N•m2/C. When the normal to the loop points in the positive y-direction the flux through the loop is Φ2 = -55 N•m2/C. When the normal to the loop points in the positive z-direction the flux though the loop is zero.

(a) Calculate the x-component of the electric field, in newtons per coulomb.
1
Expert's answer
2020-09-17T09:28:59-0400
"\Phi=EA\to E_x=\frac{\Phi_x}{A}\\E_x=\frac{\Phi_x}{\pi r^2}=\frac{65}{\pi (0.15)^2}=920\frac{N}{C}"